# Wenhao Wang

### v2

@@ 1 Introduction
The Sora [1] initializes a new era for text-to-video diffusion models, revolutionizing video generation with significant advancements. This breakthrough provides new possibilities for storytelling, immersive experiences, and content creation, as Sora [1] transforms textual descriptions into high-quality videos with ease. However, Sora [1] and other text-to-video diffusion models [2, 3, 4, 5] heavily relies on the prompts used. Despite their importance, there is no publicly available dataset focusing on text-to-video prompts, which hinders the development of these models and related researches.
@@
$\bullet$ The necessity of collecting a new prompt dataset specialized for text-to-video diffusion models. We notice that there exists a text-to-image prompt-gallery dataset, DiffusionDB [6]. By analyzing the basic information and the prompts, we conclude that the differences between our VidProM and DiffusionDB [6] mainly lies in: (1) The semantics of our prompts are significantly different from those in DiffusionDB, with our text-to-video prompts generally being more dynamic, more complex, and longer. (2) Our VidProM contains more semantically unique prompts, which are embedded by a more advanced model and collected over a longer time span. (3) We collect videos through both web scraping and local generation, whereas DiffusionDB [6] only contains web-scraped images, resulting in our dataset being more time-consuming and costly to assemble.
$\bullet$ Inspiring new research directions. The introduction of our new text-to-video prompt-gallery dataset, VidProM, opens up many exciting research directions. With the help of our VidProM, researchers can develop better, more efficient, and safer text-to-video diffusion models: (1) For better models, researchers can utilize our VidProM as a comprehensive set of prompts to evaluate their trained models, distill new models using our prompt-(generated)-video pairs, and engage in prompt engineering. (2) For more efficient models, researchers can search for related prompts in our VidProM and reconstruct new videos from similar existing videos, thereby avoiding the need to generate videos from scratch. (3) For safer models, researchers can develop specialized models to distinguish generated videos from real videos to combat misinformation, and train video copy detection models to identify potential copyright issues. Apart from diffusion models, text-video pairs can also be utilized in multimodal learning tasks such as video-text retrieval and video captioning. Our prompts and synthetic videos can help mitigate copyright concerns associated with using online videos and alleviate the difficulty of collecting video-text data.
@@ 2.2 Existing Datasets
Prompt Datasets. Existing datasets underscore the significance of compiling a set of prompts. In the text-to-text domain, studies [21] demonstrate that gathering and analyzing prompts can aid in developing language models that respond more effectively to prompts. PromptSource [22] recognizes the growing popularity of using prompts to train and query language models, and thus create a system for generating, sharing, and utilizing natural language prompts. In the text-to-image domain, DiffusionDB [6] collects a large-scale prompt-image dataset, revealing its potential to open up new avenues for research. Given the importance of prompt datasets and the new era of text-to-video generation brought by Sora [1], this paper presents the first prompt dataset specifically collected for text-to-video generation.
Figure 2: A data point in the proposed VidProM. The data point includes a prompt, a UUID, a timestamp, six NSFW probabilities, a 3072-dimensional prompt embedding, and four generated videos. Please click the corresponding links to view the complete videos: Pika, VideoCraft2, Text2Video-Zero, and ModelScope.
@@ 3 Curating VidProM
This section shows the steps of curating our VidProM, including (1) collecting source HTML files, (2) extracting and embedding prompts, (3) assigning NSFW probabilities, and (4) scraping and generating videos. In Fig. 2, we illustrate a single data point in the proposed VidProM. This data point includes a prompt, a UUID, a timestamp, six NSFW probabilities, a 3072-dimensional prompt embedding, and four generated videos. After building VidProM, we filter out prompts with high semantic similarity to each other, resulting in VidProS.
@@
Extracting and Embedding Prompts. The HTML files are then processed using regular expressions to extract prompts and time stamps. We subsequently filter out prompts used for image-to-video generation (because the images are not publicly available) and prompts without associated videos (these prompts may have been banned by Pika or hidden by the users). Finally, we remove duplicate prompts and assign a UUID to each prompt, resulting in a total of $1,672,243$ unique prompts. Because the text-to-video prompts are significantly complex and long, we use OpenAI’s text-embedding-3-large API, which supports up to $8192$ tokens, to embed all of our prompts. We retain the original 3072-dimensional output of the text-embedding-3-large model, allowing people to freely reduce their dimensionality according to their preferences using Principal Component Analysis (PCA).
Assigning NSFW Probabilities. We select the public Discord channels of Pika Labs, which prohibit NSFW content, as the source for our text-to-video prompts. Consequently, if a user submits a harmful prompt, the channel will automatically reject it. However, we find VidProM still includes NSFW prompts that were not filtered by Pika. We employ a state-of-the-art NSFW model, Detoxify [24], to assign probabilities in six aspects of NSFW content, including toxicity, obscenity, identity attack, insult, threat, and sexual explicitness, to each prompt. In Fig. 3, we visualize the number of prompts with a NSFW probability greater than $0.2$. We conclude that only a very small fraction (less than $0.5\%$) of prompts have a probability greater than $0.2$. Our dataset, VidProM, provides six separate NSFW probabilities, allowing researchers to determine a suitable threshold for filtering out potentially unsafe data specific to their tasks.
@@ 4.1 Basic Information
Images or videos. DiffusionDB focuses on images, while our VidProM is specialized in videos. Therefore, given that generating videos is much more expensive than images, it is reasonable that the number of videos in VidProM is smaller than the number of images in DiffusionDB. We also make several efforts to mitigate this disadvantage: (1) The number of source diffusion models for our VidProM is much larger than those of DiffusionDB. Our videos are generated by $4$ state-of-the-art text-to-video diffusion models, while DiffusionDB contains only images generated by Stable Diffusion. As a result, the average repetition rate per source is only $1$ for our VidProM compared to about $8.2$ for DiffusionDB. (2) We devote significantly more resources to VidProM. Unlike DiffusionDB, which only collects images through web scraping, we also deploy three open-source text-to-video models on our local servers, dedicating over $50,000$ V100 GPU hours to video generation. These efforts result in our VidProM containing more than $14$ million seconds of video.
@@ 4.2 Semantics of prompts
Takeaway: The significant difference in semantics between text-to-image prompts and our text-to-video prompts indicates the need to collect a new dataset of prompts specifically for video generation. By analyzing prompt usage distribution, future researchers can design generative models to better cater to popular prompt words and topics.
@@ 5 Inspiring New Research
The proposed million-scale VidProM dataset inspires new directions for researchers to (1) develop better, more efficient, and safer text-to-video diffusion models, and (2) advance multimodal learning from synthetic videos.
Video Generative Model Evaluation aims to assess the performance and quality of text-to-video generative models. Current evaluation efforts, such as [27, 28, 29, 30], are conducted using carefully designed and small-scale prompts. Our VidProM dataset brings imagination to this field: (1) Instead of using carefully designed prompts, researchers could consider whether their models can generalize to prompts from real users, which would make their evaluation more practical. (2) Performing evaluations on large-scale datasets will make their arguments more convincing and insightful.
@@
Text-to-Video Prompt Engineering is to optimize the interaction between humans and text-to-video models, ensuring that the models understand the task at hand and generate relevant, accurate, and coherent videos. The prompt engineering field has gained attention in large language models [34, 35], text-to-image diffusion models [36, 37], and visual in-context learning [38, 39]. However, as far as we know, there is no related research in the text-to-video community. Our VidProM provides an abundant resource for text-to-video prompt engineering. In the Appendix (Section C), we train a large language model on our VidProM for automatic text-to-video prompt completion.
@@
Fake Video Detection aims to distinguish between real videos and those generated by diffusion models. While there are some works [40, 41, 42, 43, 44, 45, 46, 47, 48, 49] focusing on fake image detection, fake video detection presents unique challenges: (1) The generalization problem: Existing fake image detectors may not generalize well to video frames (see experiments in the Appendix (Section D)). For instance, a model trained on images generated by Stable Diffusion [50] may fail to identify frames from videos generated by Pika [2] or Sora [1] as fake. (2) The efficiency problem: Currently, there is no detector that can take an entire video as input. As a result, to achieve higher accuracy, we may use fake image detectors to examine all or representative frames, which can be time-consuming. With our VidProM, researchers can (1) train specialized Fake Video Detection models on millions of generated videos, and (2) use millions of prompts to generate more videos from more diffusion models to further improve the detection performance.
Video Copy Detection for Diffusion Models aims to answer whether videos generated by diffusion models replicate the contents of existing ones (see a detailed discussion in the Appendix (Section E)). Videos generated with replicated content may infringe on the copyrights of the original videos. However, current video copy detection researches [51, 52, 53, 54] focus on detecting hand-crafted copies generated by transformations like horizontal flips and random crops, and overlook the challenge from diffusion models. Researchers can utilize our dataset VidProM as input queries and employ existing video copy detection models to pre-filter potential replications. Subsequently, these filtered videos can be manually annotated to determine whether they are indeed replications and regarded as training sources.
Multimodal Learning from Synthetic Videos. Multimodal learning related to text-video interactions include tasks such as video-text retrieval [55, 56, 57] and video captioning [58, 59, 60]. They are typically trained on large-scale video datasets collected from online sources such as YouTube and Shutterstock. However, there are two main concerns: (1) copyright issues, for example, the WebVid-10M dataset [16] is no longer publicly available due to its infringement on Shutterstock’s copyright; and (2) the difficulty in collecting high-quality video-text data, as manual labeling is time-consuming. Our VidProM offers a promising research direction, i.e. learning from synthetic videos, which are free from copyright issues and do not require any manual labeling.
@@ 6 Conclusion
Limitation. We recognize that the videos currently generated are short and not of the highest quality. In the future, we intend to enhance our dataset by incorporating high-quality videos produced by more advanced models, like Sora, using our long and detailed prompts.
@@ Appendix B Domain Gap between Video Captions and Text-to-Video Prompts
Appendix C Automatic Text-to-video Prompt Completion
Appendix D Benchmarking Existing Fake Image Detection on VidProM
@@
Appendix E Video Copy Detection for Diffusion Models
@@ E.2 Observations
E.3 Limitations and Future Directions
@@
Appendix F Author Statement
We, the undersigned, hereby affirm that we bear full responsibility for the content provided herein. We guarantee that all data and materials used and presented are compliant with applicable laws and are free from any encroachments on intellectual property rights or any other legal rights of third parties. In the event of any violations, we accept full liability and will address all associated legal consequences.
Appendix G Hosting, licensing, and maintenance plan.
The data is hosted by Hugging Face, Inc. at https://huggingface.co/datasets/WenhaoWang/VidProM. The prompts and videos generated by Pika in our VidProM are licensed under the CC BY-NC 4.0 license from the Pika Discord, allowing for non-commercial use with attribution. Additionally, similar to their original repositories, the videos from VideoCraft2, Text2Video-Zero, and ModelScope are released under the Apache license, the CreativeML Open RAIL-M license, and the CC BY-NC 4.0 license, respectively. We commit to maintaining the dataset, which includes regular updates and corrections as necessary.
Appendix H Data Sheet for VidProM
@@
The dataset was created by Wenhao Wang (University of Technology Sydney), Yifan Sun (Baidu Inc.), and Yi Yang (Zhejiang University).
@@
All messages were generated between July 2023 and February 2023. VidProM includes the generation timestamps of all videos.
@@
All tasks that utilize this dataset should follow the licensing policies and the regulations posed by Pika Lab.
@@
Yes, we will monitor cases when users can report harmful images and creators can remove their videos. We will update the dataset bimonthly. Updates will be posted on the project website https://vidprom.github.io/. In the future, we intend to enhance our dataset by incorporating high-quality videos produced by more advanced models, like Sora, using our long and detailed prompts.

### v4

@@ 1 Introduction
The Sora [1] initializes a new era for text-to-video diffusion models, revolutionizing video generation with significant advancements. This breakthrough provides new possibilities for storytelling, immersive experiences, and content creation, as Sora [1] transforms textual descriptions into high-quality videos with ease. However, Sora [1] and other text-to-video diffusion models [2, 3, 4, 5] heavily relies on the prompts used. Despite their importance, there is no publicly available dataset focusing on text-to-video prompts, which may hinder the development of these models and related researches.
@@
$\bullet$ The necessity of collecting a new prompt dataset specialized for text-to-video diffusion models. We notice that there exists a text-to-image prompt-gallery dataset, DiffusionDB [6]. By analyzing the basic information and the prompts, we conclude that the differences between our VidProM and DiffusionDB [6] mainly lies in: (1) Semantics: The semantics of our prompts are significantly different from those in DiffusionDB, with our text-to-video prompts generally being more dynamic, more complex, and longer. (2) Modality: DiffusionDB [6] focuses on images, while our VidProM is specialized in videos. (3) Techniques: We utilize some recent advanced techniques, such as latest text embedding model (OpenAI-text-embedding-3-large), to build our VidProM.
$\bullet$ Inspiring new research directions. The introduction of our new text-to-video prompt-gallery dataset, VidProM, opens up many exciting research directions. With the help of our VidProM, researchers can develop better, more efficient, and safer text-to-video diffusion models: (1) For better models, researchers can utilize our VidProM as a comprehensive set of prompts to evaluate their trained models, distill new models using our prompt-(generated)-video pairs, and engage in prompt engineering. (2) For more efficient models, researchers can search for related prompts in our VidProM and reconstruct new videos from similar existing videos, thereby avoiding the need to generate videos from scratch. (3) For safer models, researchers can develop specialized models to distinguish generated videos from real videos to combat misinformation, and train video copy detection models to identify potential copyright issues.
@@
Figure 2: A data point in the proposed VidProM. Please click the corresponding links to view the complete videos: Pika, VideoCraft2, Text2Video-Zero, and ModelScope. To better understand our dataset, yon can also download $10,000$ randomly-selected data points from here.
@@ 2.2 Existing Datasets
Prompt Datasets. Existing datasets underscore the significance of compiling a set of prompts. In the text-to-text domain, studies [21] demonstrate that gathering and analyzing prompts can aid in developing language models that respond more effectively to prompts. PromptSource [22] recognizes the growing popularity of using prompts to train and query language models, and thus create a system for generating, sharing, and utilizing natural language prompts. In the text-to-image domain, DiffusionDB [6], which is nominated for the best paper of ACL 2023, collects a large-scale prompt-image dataset, revealing its potential to open up new avenues for research. Given the importance of prompt datasets and the new era of text-to-video generation brought by Sora [1], this paper presents the first prompt dataset specifically collected for text-to-video generation.
@@ 3 Curating VidProM
In Fig. 2, we illustrate a single data point in the proposed VidProM. This data point includes a prompt, a UUID, a timestamp, six NSFW probabilities, a 3072-dimensional prompt embedding, and four generated videos. We show the steps of curating our VidProM in this section.
@@
Extracting and Embedding Prompts. The HTML files are then processed using regular expressions to extract prompts and time stamps. We subsequently filter out prompts used for image-to-video generation (because the images are not publicly available) and prompts without associated videos (these prompts may have been banned by Pika or hidden by the users). Finally, we remove duplicate prompts and assign a UUID to each prompt, resulting in a total of $1,672,243$ unique prompts. Because the text-to-video prompts are significantly complex and long, we use OpenAI’s text-embedding-3-large API, which supports up to $8192$ tokens, to embed all of our prompts. We retain the original 3072-dimensional output, allowing any customized dimensionality reduction.
Assigning NSFW Probabilities. We select the public Discord channels of Pika Labs, which prohibit NSFW content, as the source for our text-to-video prompts. Consequently, if a user submits a harmful prompt, the channel will automatically reject it. However, we find VidProM still includes NSFW prompts that were not filtered by Pika. We employ a state-of-the-art NSFW model, Detoxify [24], to assign probabilities in six aspects of NSFW content, including toxicity, obscenity, identity attack, insult, threat, and sexual explicitness, to each prompt. In Fig. 3, we visualize the number of prompts with a NSFW probability greater than $0.2$. We conclude that only a very small fraction (less than $0.5\%$) of prompts have a probability greater than $0.2$. We provide six separate NSFW probabilities, enabling researchers to set a suitable threshold for filtering out potentially unsafe data for their tasks.
@@ 4.1 Basic Information
Images or videos. DiffusionDB focuses on images, while our VidProM is specialized in videos. Therefore, given that generating videos is much more expensive than images, it is reasonable that the number of videos in VidProM is smaller than the number of images in DiffusionDB. We also make several efforts to mitigate this disadvantage: (1) The number of source diffusion models for our VidProM is much larger than those of DiffusionDB. Our videos are generated by $4$ state-of-the-art text-to-video diffusion models, while DiffusionDB contains only images generated by Stable Diffusion. As a result, the average repetition rate per source is only $1$ for our VidProM compared to about $8.2$ for DiffusionDB. (2) We devote significantly more resources to VidProM. Unlike DiffusionDB, which only collects images through web scraping, we also deploy three open-source text-to-video models on our local servers, dedicating over $50,000$ V100 GPU hours to video generation. These efforts result in more than $14$ million seconds of video. In the future, researchers can also generate more videos using the prompts with more advanced text-to-video diffusion models.
@@ 4.2 Semantics of prompts
Takeaway: The significant difference in semantics between text-to-image prompts and our text-to-video prompts indicates the need to collect a new dataset of prompts specifically for video generation. By analyzing prompt usage distribution, future researchers can design generative models to better cater to popular topics in prompts.
@@ 5 Inspiring New Research
The proposed million-scale VidProM dataset inspires new directions for researchers to develop better, more efficient, and safer text-to-video diffusion models.
Video Generative Model Evaluation aims to assess the performance and quality of text-to-video generative models. Current evaluation efforts, such as [27, 28, 29, 30], are conducted using carefully designed and small-scale prompts. Our VidProM dataset brings imagination to this field: (1) Instead of using carefully designed prompts, researchers could consider whether their models can generalize to prompts from real users, which would make their evaluation more practical. (2) Performing evaluations on large-scale datasets will make their arguments more convincing.
@@
Text-to-Video Prompt Engineering is to optimize the interaction between humans and text-to-video models, ensuring that the models understand the task at hand and generate relevant, accurate, and coherent videos. The prompt engineering field has gained attention in large language models [34, 35], text-to-image diffusion models [36, 37], and visual in-context learning [38, 39]. However, as far as we know, there is no related research in the text-to-video community. Our VidProM provides an abundant resource for text-to-video prompt engineering. In the Section 6, we train a large language model on our VidProM for automatic text-to-video prompt completion for instance.
@@
Fake Video Detection aims to distinguish between real videos and those generated by diffusion models. While there are some works [40, 41, 42, 43, 44, 45, 46, 47, 48, 49] focusing on fake image detection, fake video detection presents unique challenges: (1) The generalization problem: Existing fake image detectors may not generalize well to video frames (see experiments in the Appendix (Section C)). For instance, a model trained on images generated by Stable Diffusion [50] may fail to identify frames from videos generated by Pika [2] or Sora [1] as fake. (2) The efficiency problem: Currently, there is no detector that can take an entire video as input. As a result, to achieve higher accuracy, we may use fake image detectors to examine all or representative frames, which can be time-consuming. With our VidProM, researchers can (1) train specialized Fake Video Detection models on millions of generated videos, and (2) use millions of prompts to generate more videos from more diffusion models to further improve the detection performance.
Video Copy Detection for Diffusion Models aims to answer whether videos generated by diffusion models replicate the contents of existing ones (see a detailed discussion in the Appendix (Section D)). Videos generated with replicated content may infringe on the copyrights of the original videos. However, current video copy detection researches [51, 52, 53, 54] focus on detecting hand-crafted copies generated by transformations like horizontal flips and random crops, and overlook the challenge from diffusion models. Researchers can utilize our dataset VidProM as input queries and employ existing video copy detection models to pre-filter potential replications. Subsequently, these filtered videos can be manually annotated to determine whether they are indeed replications and regarded as training sources.
6 Automatic Text-to-video Prompt Completion
@@ 7 Conclusion
Limitation. We recognize that the videos currently generated are short and not of the highest quality. In the future, we intend to enhance our dataset by incorporating high-quality videos produced by more advanced models, like Sora, using our long and detailed prompts. At the time of camera-ready version, we utilize three new powerful text-to-video models models to generate 10,000 videos with each for example: 10,000 videos of 8 seconds at 720p quality for StreamingT2V [58], 10,000 videos of 8 seconds at 720p quality for Open-Sora 1.2 [59], and 10,000 videos of 6 seconds at 720p quality for CogVideoX-2B [60]. They are publicly available at https://huggingface.co/datasets/WenhaoWang/VidProM/tree/main/example.
@@ Appendix B Domain Gap between Video Captions and Text-to-Video Prompts
Appendix C Benchmarking Existing Fake Image Detection on VidProM
@@
Appendix D Video Copy Detection for Diffusion Models
@@ D.2 Observations
D.3 Limitations and Future Directions
@@
Appendix E Data Sheet for VidProM
@@
The dataset was created by Wenhao Wang (University of Technology Sydney) and Yi Yang (Zhejiang University).
@@
All messages were generated between July 2023 and February 2024. VidProM includes the generation timestamps of all videos.
@@
All tasks that utilize this dataset should follow the licensing policies posed by Pika Lab.
@@
Yes, we will monitor cases when users can report harmful videos and creators can remove their videos. We will update the dataset bimonthly. Updates will be posted on the project website https://vidprom.github.io/. In the future, we intend to enhance our dataset by incorporating high-quality videos produced by more advanced models, like Sora, using our detailed prompts.